OpenStudy (anonymous):

|x+3|+8=6 Help me solve this problem

OpenStudy (anonymous):

x+3+8=6 x=6-8-3 x=-5

OpenStudy (anonymous):

x=-5

OpenStudy (anonymous):

technically that doesn't work if you put -5 back in...it doesn't work

OpenStudy (mathhelp346):

x=-15?

OpenStudy (mathhelp346):

no wait

OpenStudy (anonymous):

|(-5)+3|+8=6 |-2|+8=6 2+8=6 10=6

OpenStudy (anonymous):

but those solutions are extraneous...

OpenStudy (anonymous):

You get |x+3| = -2 but the absolute value is never negative, so the equation has no solutions.

OpenStudy (mathhelp346):

wait is x = -5

OpenStudy (anonymous):

yea that's what i'm trying to say :) it has solutions but those solutions doesn't work

OpenStudy (mathhelp346):

-5+3=-2 -2+8= 6

OpenStudy (anonymous):

Which means it doesn't have solutions :D.

OpenStudy (mathhelp346):

hmm?

OpenStudy (mathhelp346):

?

OpenStudy (anonymous):

because it has absolute value...you can't just take away the absolute value

OpenStudy (anonymous):

Your right the answer is no solution , I was double checking my work to make sure that I was understanding the problem correctly. Thanks for the help.

OpenStudy (mathhelp346):

medal? :D